# Spline-FRIDA: Towards Diverse, Humanlike Robot Painting Styles with a Sample-Efficient, Differentiable Brush Stroke Model

### v1

@@ II Introduction
Paintings and drawings are used to convey messages of emotion, cultural values, and shared experiences. While these aspects can be conveyed by the objects or subjects within the painting, style is perhaps just as important to expressing those messages.
“The ‘same’ content represented in a different form—in a different medium or mode or style or language—is not the same: what is the same through all variations of the form is only a tenuous abstraction, a précis of the full content” [1]
In the visual art space, there is evidence that people find the style of an image to be even more crucial than its content when interpreting the meaning of a generated image [2]. In particular, patterns in the shapes of individual strokes within a painting can contribute to the overall style and aesthetic of an artwork. Two examples of this can be seen in Figure 1. The drawing on the left uses a style with long, jagged strokes, giving the parrot a fluffy texture. On the other hand, the drawing on the right is composed of small circular strokes, resulting in a different texture. In both cases, the stroke shapes are crucial for defining the painting’s style and therefore the expression of the message that the artist intends to convey.
@@
In this paper, we explore how stroke-level style control can be added to FRIDA [9], an open-source framework for robot painting. FRIDA uses simple Bézier curves as stroke primitives, but this representation lacks human style. To adapt FRIDA to generate more human-like brush strokes, we use motion capture to record human drawings with real-world brushes and markers on paper. We model these recorded trajectories with an autoencoder. In addition, we find that FRIDA’s brush stroke dynamics model (responsible for converting stroke parameters into rendered images) is unable to model the complex strokes that humans create because it is a purely data-driven approach that lacks priors, leading to a large Sim2Real gap. Thus, we introduce a novel brush stroke dynamics model, Traj2Stroke, that is more capable of modeling the collected human stroke data. Traj2Stroke is on par with FRIDA’s dynamics model for simple strokes, and is much better for complex strokes.
@@
  • We present a successful approach to modeling styles of brushstroke trajectories using variational autoencoders, TrajVAE. This gives users control over the shapes of individual strokes used for a robot painting.
  • We demonstrate the viability of using motion capture to track a human artist’s drawing in real time. We are able to accurately recover stroke trajectories from the raw position data. We open source a small dataset of human drawings collected with this method along with the entirety of the Spline-FRIDA code.
@@ III Related Work
Stroke-Based Rendering (SBR) involves arranging primitive shapes to create an image, often with the goal of replicating some target image. Many recent works use forward prediction methods, in which a neural network learns to output the next stroke to add [10, 11, 12], as well as optimization-based methods, where stroke parameters are passed through a differentiable rendering pipeline and optimized via backpropagation [9, 13, 14].
@@ III-A Stroke Primitives
Most SBR research is focused on global planning and propose new algorithms to arrange stroke primitives. On the other hand, there has been little research into how the stroke primitives themselves should be defined. Some works use definitions that would be difficult to replicate on a physical robot. For instance, Learning to Paint defines strokes as translucent Bézier curves with arbitrary thicknesses [11]. Schaldenbrand et al. found that when Learning to Paint’s neural network was restricted to outputting realistic brush strokes, the quality of generated images suffered [10]. Paint Transformer uses a mask of a brush stroke that can be transformed, resized, and recolored [12]. However, the wide range of stroke sizes presents a problem, as some strokes can be half the size of the canvas, while others are thin enough to capture fine details. In practice, implementing a setup in real life that allows the robot arm to make arbitrarily sized brush strokes is infeasible.
@@ IV Background
Our work is based on FRIDA [9], a robotic system that can paint what users describe via text prompts, images, or even audio recordings [21] in an interactive and collaborative manner [22].
FRIDA is an optimization-based SBR method. It works by using a fully differentiable rendering module that takes in a list of stroke parameters and outputs a rendered drawing. Initially, the planner assigns random values to the stroke parameters. It optimizes these values by passing them through the rendering module, computing a loss according to the user-specified objectives, and back-propagating it to the parameters. After many iterations, the system converges on a set of stroke parameters that can be executed on the robot.
@@
FRIDA renders strokes using a convolutional neural network (CNN). Although the CNN approach works well for short, simple strokes made with a paintbrush, it has trouble converging when longer, thinner Sharpie marker strokes are used, as illustrated in Figure 2. We believe this issue arises because thin ground truth strokes require the renderer to achieve a higher level of precision, complicating its task. Consequently, the CNN often produces blotchy, incomplete predictions. This is problematic since the errors compound during the planning stage (illustrated on the right side of Figure 2).
Another avenue for improvement is FRIDA’s stroke representation. FRIDA uses quadratic Bézier curves defined by three parameters—length, height above the canvas, and bend—to represent strokes. While this straightforward approach offers convenience, it falls short in capturing the complexity and variety inherent in human drawings. In this context, we propose a novel stroke representation and a new renderer to model long, complex trajectories that can enable diverse styles.
@@ V-A Overview
Our approach to stroke modeling and renderering consists of (1) capturing and processing human demonstration data using motion capture technology, (2) modeling these trajectories by training an autoencoder, TrajVAE, and (3) using Real2Sim2Real methodology to fine-tune our novel rendering approach, Traj2Stroke. Importantly, we switch from FRIDA’s quadratic Bézier curve representation to a new representation in which trajectories can be arbitrary polyline splines.
Figure 3: Spline-FRIDA Brush Stroke Dynamics Model. The inputs are a latent vector $z$ and an offset $\Delta$. $z$ is fed through the decoder of a TrajVAE, generating a raw trajectory, which is then rotated and translated according to $\Delta$. We then process the trajectory segments independently, obtaining darkness values for each. Finally, we take the max darkness over all segments.
A visualization of the rendering pipeline for a single stroke is shown in Fig. 3. Stroke parameters are expressed as $(z,\Delta)$ pairs, where $z$ is the latent vector describing the shape of the stroke, and $\Delta$ represents its rotation and translation on the canvas. $z$ is fed into the decoder of a variational autoencoder, which outputs a trajectory. This is then rotated and translated according to $\Delta$. The next part of the pipeline is the Traj2Stroke model, which differentiably renders the trajectory using a sequence of tensor operations. The output is an array of darkness values in the interval $[0,1]$.
Finally, the individually rendered strokes are stamped onto the same canvas to create the final painting. Importantly, each step in this process is differentiable, which allows gradients to flow from the painting back to the stroke parameters.
@@ V-B Motion Capture Drawing Recording and Processing
A motion capture system is utilized instead of a drawing tablet to capture human brushstroke trajectories, as it better replicates the robot’s environment (both humans and robots use identical physical tools).
@@
While the artist sketches, we continuously track the positions and orientations of the canvas and pen. For each frame, we calculate the position of the pen tip and determine its distance from the canvas. If this value is below a threshold, we consider the pen to be in contact with the paper. Consecutive positions where this is the case are merged into trajectories. Each trajectory is then standardized by translating it to the origin and rotating it to be horizontal (ending at $y=0$), as seen in Figure 4. We also resample each trajectory to have exactly $32$ points.
Thus, each human brushstroke trajectory is modeled as a polyline (piecewise linear) going through $32$ control points. This polyline is encoded as a $32\times 3$ tensor. The coordinates $(x,y,h)$ of each control point are defined by $x$ and $y$ as horizontal displacements (in the plane of the canvas) and $h$ as vertical displacement (elevation of the brush above the canvas).
@@ V-C TrajVAE
After collecting and processing the motion capture data, we train variational autoencoders [23] to model these stroke trajectories. We name these TrajVAEs. During training, a TrajVAE takes a trajectory as input, passes it through an encoder that compresses it to a latent vector of size $64$, and then sends it through a decoder to turn it back into a trajectory. We minimize the mean squared error between the input and output trajectories.
During a single motion capture recording, we record roughly 100 human-drawn trajectories. We found that this is not enough data for a TrajVAE to converge. Instead, we pretrain each TrajVAE on trajectories aggregated from multiple recording sessions, then fine-tune it on a single session to capture a more specific style. Each model converges very fast (less than a minute) and only requires a few (<20) trajectories in the fine-tuning dataset.
@@
Our motion capture setup is not precise enough to capture the variation in the vertical component, so the TrajVAE does not model the $h$ values of the stroke. Instead, four values of $h$ are used as parameters which are linearly interpolated into $n$ values and stacked with the $x$ and $y$ trajectory values.
@@ V-D Traj2Stroke Model
The TrajVAE model outputs a trajectory that the robot should draw, but to predict the appearance of the stroke given this trajectory, we developed a novel rendering approach that we call Traj2Stroke. Traj2Stroke takes a trajectory, as well as positional and rotational offsets, and renders it as an $H\times W$ image. Importantly, this process is differentiable so that the planning process can backpropagate through it.
To train this model, we randomly sample trajectories from TrajVAE, execute them on the robot, and take before/after pictures of the canvas for each stroke. Next, we input the sampled trajectories into the Traj2Stroke model to get predicted stroke masks. These masks are stamped onto the before-stroke pictures, and the resulting prediction is compared to the after-stroke pictures. We minimize a weighted L1 loss that places higher weight on pixels covered by the new stroke. A separate Traj2Stroke model must be trained for each drawing medium (marker/brush).
@@
We proceed by rendering each its $n-1$ segments separately. Fix an arbitrary $k$, and note that segment $k$ goes from $(x_{k}^{\prime},y_{k}^{\prime},h_{k})$ to $(x_{k+1}^{\prime},y_{k+1}^{\prime},h_{k+1})$.
@@
In total, the Traj2Stroke model has only $7$ learnable parameters:
$(x_{m},y_{m},x_{b},y_{b},\alpha,\beta,c).$
Essentially, the output is constrained to be a polyline that follows the trajectory of the brush. During training, the model only learns the small affine transformation, the thickness parameters, and the darkness dropoff. As we will show in Section VI-C, this design is more precise and more robust to overfitting compared to CNN renderers.
@@ VI Results
Figure 5 shows an array of drawings produced by Spline-FRIDA. We capture five human drawings using the mocap system, each with distinct stroke styles, which are presented in the top row. Each human drawing is used to fine-tune a separate TrajVAE, resulting in five unique TrajVAEs. Each TrajVAE is then used to plan a series of drawings with various objectives. These objectives are displayed in the left column.
@@
Figure 5: Example drawings made by Spline-FRIDA. Each column represents a distinct trajectory style and each row uses a different objective. For these examples, we maximize the CLIP similarity score between the canvas and the objective. The top row contains original drawings made by human artists on our mocap system. One VAE was fine-tuned on each human drawing and used to plan the drawings in each column.
@@ VI-A Human Evaluations
To what extent is Spline-FRIDA able to capture the stroke style of a drawing? And, in general, are Spline-FRIDA’s drawings better than those made by FRIDA? To obtain quantitative answers to these questions, we conducted a survey and released it to 100 participants on Amazon Mechanical Turk.
For the first part of the survey, we asked participants to match Spline-FRIDA drawings with human drawings that have the same stroke style. More specifically, for each participant, we pick a random row of Figure 5 (technically, an extended version with 11 different objectives), shuffle it, and for each drawing, ask them to pick the closest human drawing in stroke style. We told participants to “focus on the characteristics of individual strokes, such as their trajectories, shapes, and curves.” The results of this experiment are seen in Figure 6.
The high values along the diagonal suggest that, in general, participants were able to choose the correct human drawing used to style each robot drawing. Style 5, which corresponds to the drawings made with small circles, seemed to be particularly distinguishable. Meanwhile, style 1 (which uses very short strokes) was often confused with style 4 (which also uses short strokes but is more curly), and style 2 (which uses long strokes) was confused with style 3 (which also can have long strokes, but tend to be completely straight). Nevertheless, all five encoded styles are most strongly associated with the correct human drawings.
@@
TABLE I: Opinions on FRIDA vs Spline-FRIDA. Each cell shows the number of participants that chose the system for the given question. Overall, participants thought that compared to FRIDA, drawings made by Spline-FRIDA were more human-like, higher quality, more true to the objective, and more artistic.
@@ VI-B Trajectory Distributions
In figure 7, we visualize the latent space for our TrajVAE trained on multiple drawing sessions. We encode all of the human trajectories into latent vectors, then project them down to 2 dimensions using t-SNE [24]. We then draw each human trajectory at its corresponding 2d coordinates.
We observe human trajectories spread throughout the space, forming several homogeneous clusters. This structured organization indicates that TrajVAE effectively learns a correlation between trajectories and latent vectors. Consequently, an optimization-based planning algorithm is likely to be effective.
@@ VI-C Brush Stroke Dynamics Modeling Experiments
The purpose of a stroke dyanmics model is to differentiably render trajectories. We experiment with a variety of methods to do this and evaluate them both quantitatively and qualitatively in a controlled experiment. In total, we evaluated four methods:
@@ VI-C1 CNN
Our baseline, a convolutional neural network. This is analogous to FRIDA’s renderer architecture, except it takes in full trajectory rather than the parameterization (length, bend, height) that FRIDA uses.

### v2

@@ II Introduction
Paintings and drawings are used to convey messages of emotion, cultural values, and shared experiences. While these aspects can be conveyed by the objects or subjects within the painting, style is perhaps just as important to expressing those messages [1, 2]. In particular, patterns in the shapes of individual strokes within a painting can contribute to the overall style and aesthetic of an artwork. Some examples can be seen in Figure 1, with the left drawings using long strokes and the right drawings using small, circular ones. In both cases, the stroke shapes are crucial for defining the painting’s style and therefore the expression of the message that the artist intends to convey.
@@
Prior work has mostly focused on planning paintings using basic stroke representations such as Bézier curves [9, 10] and only fixating on the style of the overall image [2, 11]. In this paper, we focus on how intra-stroke style control can be implemented.
Our work uses motion capture to record human drawings with real-world brushes and markers on paper. We model these recorded trajectories with an autoencoder, TrajVAE. We also introduce a novel brush stroke dynamics model, Traj2Stroke, which predicts the 2d outline of stroke given its trajectory.
@@
  • We present a successful approach to modeling human styles of brushstroke trajectories using variational autoencoders.
@@ III Related Work
Stroke-Based Rendering (SBR) involves arranging primitive shapes to create an image, often with the goal of replicating some target image. Some recent works use forward prediction methods, in which a neural network learns to output the next stroke to add [12, 9, 13], while others use optimization-based methods, where stroke parameters are passed through a differentiable rendering pipeline and optimized via backpropagation [10, 11, 14].
@@ III-A Stroke Primitives
Most SBR research is focused on global planning and propose new algorithms to arrange stroke primitives. On the other hand, there has been little research into how the stroke primitives themselves should be defined. Some works use definitions that would be difficult to replicate on a physical robot. For instance, Learning to Paint defines strokes as translucent Bézier curves with arbitrary thicknesses [9]. Schaldenbrand et al. found that when their system was restricted to outputting more realistic brush strokes by making them opaque and limiting the sizes, the quality of generated images suffered [12]. Paint Transformer uses a mask of a brush stroke that can be transformed, resized, and recolored [13]. This arbitrary sizing of strokes without loss of precision would be very difficult to implement in hardware.
@@ IV Background
Our work is built on FRIDA [10], a robotic system that paints what users describe via text prompts, images, or audio recordings [21] in an interactive and collaborative manner [22].
@@
FRIDA renders strokes using a convolutional neural network (CNN). Although the CNN works well for short, simple strokes made with a paintbrush, it has trouble converging when longer, thinner Sharpie marker strokes are used, as illustrated in Figure 2. We believe this issue arises because thin ground truth strokes require the renderer to achieve a higher level of precision, complicating its task. Consequently, the CNN often produces blotchy, incomplete predictions. This is problematic since the errors compound during the planning stage (illustrated on the right side of Figure 2).
Additionally, FRIDA’s stroke representation is limited. FRIDA, like many other computer painting works [2, 9, 11], uses quadratic Bézier curves defined by three parameters—length, height above the canvas, and bend—to represent strokes. While this straightforward approach offers convenience, it falls short in capturing the complexity and variety inherent in human drawings. In this context, we propose a novel stroke representation and a new renderer to model long, complex trajectories that can enable diverse styles.
@@ V-A Overview
Our approach to stroke modeling and rendering consists of (1) capturing and processing human demonstration data using motion capture technology, (2) modeling these trajectories by training an autoencoder, TrajVAE, (3) using Real2Sim2Real methodology to fine-tune our novel rendering approach, Traj2Stroke, and (4) planning using gradient descent to optimize a set of brush stroke parameters through our dynamics model to decrease the feature-space loss between a given image and the predicted painting.
Figure 3: Traj2Stroke. The inputs are a latent vector $z$ and an offset $\Delta$. $z$ is fed through the decoder of a TrajVAE, generating a raw trajectory, which is then rotated and translated according to $\Delta$. We then process the trajectory segments independently, obtaining darkness values for each. Finally, we take the max darkness over all segments.
@@ V-B Motion Capture Drawing Recording and Processing
We utilize a motion capture system consisting of OptiTrack cameras and Motive software to capture human brushstroke trajectories.
@@
While the artist sketches, we continuously track the positions and orientations of the canvas and pen. Using a manual measurement of the length of the pen, we are able to calculate the position of the pen tip and determine its distance from the canvas. If this value is below a threshold, we consider the pen to be in contact with the paper. Consecutive positions where this is the case are merged into trajectories. Each trajectory is then standardized by translating it to the origin and rotating it to be horizontal (ending at $y=0$), as seen in Figure 4, to reduce variation for sample-efficient modeling. It is worth noting that this normalization has a tradeoff: it assumes trajectory style is not affected by position/rotation on the canvas. We also resample each trajectory to have exactly $n$ points (in practice $n=32$).
Thus, each human brushstroke trajectory is modeled as a polyline (piecewise linear) going through $n$ control points. This polyline is encoded as a $n\times 3$ tensor. The coordinates $(x,y,h)$ of each control point are defined by $x$ and $y$ as horizontal displacements (in the plane of the canvas) and $h$ as vertical displacement (elevation of the brush above the canvas).
@@ V-C TrajVAE
After collecting and processing the motion capture data, we train variational autoencoders [23] to model these stroke trajectories. We name these TrajVAEs. During training, a TrajVAE takes a trajectory as input, passes it through an MLP encoder that compresses it to a latent vector of size $64$, and then sends it through a MLP decoder to turn it back into a trajectory. We minimize the mean squared error between the input and output trajectories.
We typically record between 20 and 200 human-drawn trajectories per drawing, but found that this is not enough data to robustly train a TrajVAE from scratch. Instead, we pretrain each TrajVAE on trajectories aggregated from multiple recording sessions, then fine-tune it on a single session to capture a more specific style. Each model converges very fast (less than a minute) and only requires a few (<20) trajectories in the fine-tuning dataset.
@@
Our motion capture device struggles to capture the vertical position of the drawing utensil’s tip with enough precision. This is because a small height difference can drastically affect the thickness of a stroke. Thus, rather than explicitly modeling the height with TrajVAE, in practice we optimize it as separate stroke parameters during the planning process.
Figure 5: Planning a Painting. As described in Section V-F, Spline-FRIDA plans a painting by optimizing the brush stroke parameters through the dynamics model to decrease a features space loss between a given image and the planned painting. Whereas FRIDA models brush strokes as simple Bézier curves, Spline-FRIDA uses trajectories which enable highly flexible brush strokes.
@@ V-D Traj2Stroke Model
The TrajVAE model outputs a trajectory that the robot should draw, but to predict the appearance of the stroke given this trajectory, we developed a novel rendering approach that we call Traj2Stroke. Traj2Stroke takes a trajectory, as well as positional and rotational offsets, and renders it as a grayscale image with height and width dimensions, $H\times W$. Importantly, this process is differentiable so that the planning process can backpropagate through it.
To train this model, we randomly sample trajectories from TrajVAE, execute them on the robot, and take before/after pictures of the canvas for each stroke. Next, we input the sampled trajectories into the Traj2Stroke model to get predicted stroke masks. These masks are stamped onto the before-stroke pictures, and the resulting prediction is compared to the after-stroke pictures. We minimize a weighted L1 loss that places higher weight on pixels covered by the new stroke. In practice, collecting the dataset (including setup and execution) takes around an hour, and training takes around 20 minutes to converge on our single-GPU system.
A separate Traj2Stroke model must be trained for each drawing medium (marker/brush), but is robust to out-of-distribution trajectories. This means that when we obtain a new TrajVAE, we can almost always plug-and-play it into the system without needing to collect new data and retraining Traj2Stroke. This is very convenient, as collecting the Traj2Stroke dataset is usually the most time-consuming part of preparations.
@@
We proceed by rendering each of its $n-1$ segments separately. Fix an arbitrary $k$, and note that segment $k$ goes from $(x_{k}^{\prime},y_{k}^{\prime},h_{k})$ to $(x_{k+1}^{\prime},y_{k+1}^{\prime},h_{k+1})$.
@@
In total, the Traj2Stroke model has only $7$ learnable parameters: $m_{x}$, $m_{y}$, $b_{x}$, $b_{y}$, $\alpha$, $\beta$, and $c$.
V-E Stroke Composition
We define each brush stroke action as a set of parameters: TrajVAE latent vector $z$, pose offsets $\Delta$, and RGB color $\rho$. Given $z$ and $\Delta$Figure 3 illustrates how Spline-FRIDA predicts the shape of a single stroke. Next, the rendered stroke is colorized by duplicating it to 3-channels and multiplying each channel based on the stroke color as seen in Figure 5. We can predict how this stroke $s$ will appear once it is performed on a canvas $c_{t}$ by stamping it via an alpha blending formula.
V-F Painting and Drawing Planning
To plan a painting or drawing, we follow the FRIDA [10] planning algorithm which plans paintings using an optimization loop, depicted in Figure 5. A user-specified number of brush stroke actions are randomly initialized. At each optimization step, the current canvas is compared to the user-specified target image forming a loss value. In practice, features from the planned painting and target image are extracted using pretrained neural networks (e.g., CLIP [24]) and compared using cosine similarity as introduced in [25]. The loss is back-propagated through the dynamics model to the brush stroke parameters which are updated using gradient descent.
If the robot is painting in color, the color parameters are optimized as continuous RGB values during initial iterations. In the last 10% of optimization iterations, the algorithm discretizes the colors to a user-specified number using K-Means clustering. After optimizing for 2000 iterations, the system shows which colors of paint need to be mixed in a graphical user-interface. The user mixes these paint colors, provides them to the robot, then the robot can begin painting.
@@ VI Results
Figure 6: Example drawings made by Spline-FRIDA. Each column represents a distinct trajectory style and each row uses a different objective. The top row contains original drawings made by human artists on our mocap system. One VAE was fine-tuned on each human drawing and used to plan the drawings in each column.
Figure 6 shows an array of drawings produced by Spline-FRIDA. We hand-pick five human drawings from members of our lab using the mocap system, each with distinct stroke styles, which are presented in the top row. Each human drawing is used to fine-tune a separate TrajVAE, resulting in five unique TrajVAEs. Each TrajVAE is then used to plan a series of drawings with various objectives. These objectives are displayed in the left column.
@@ VI-A Human Evaluations
To what extent is Spline-FRIDA able to capture the stroke style of a drawing? And, in general, are Spline-FRIDA’s drawings better than those made by FRIDA? These questions are subjective and difficult to answer with automatic metrics. To obtain quantitative results, we conducted a survey and released it to 100 participants on Amazon Mechanical Turk.
For the first part of the survey, we asked participants to match Spline-FRIDA drawings with human drawings that have the same stroke style. More specifically, for each participant, we selected a random human drawing, along with five robot drawings (a random row of Figure 6), and asked them to pick the robot drawing that best matched the style of the human drawing. We told participants to “focus on the characteristics of individual strokes, such as their trajectories, shapes, and curves.” The results of this experiment are seen in Figure 7.
The high values of Figure 7 along the diagonal suggest that, in general, participants were able to choose the correct human drawing used to style each robot drawing. Style 5 seemed to be particularly distinguishable. Meanwhile, style 1 was often confused with style 4, and style 2 was confused with style 3. Nevertheless, all five encoded styles are most strongly associated with the correct human drawings.
@@
TABLE I: Opinions on FRIDA vs Spline-FRIDA. Each cell shows the number of participants that chose the system for the given question. Overall, participants thought that compared to the Bèzier curve representation of FRIDA, drawings made by Spline-FRIDA were more human-like, higher quality, more true to the objective, and more artistic.
@@ VI-B Trajectory Distributions
In Figure 8, we visualize the latent space for our TrajVAE trained on multiple drawing sessions. We encode all of the human trajectories into latent vectors, then project them down to 2 dimensions using t-SNE [26]. We then draw each human trajectory at its corresponding 2d coordinates. We observe human trajectories spread throughout the space, forming several homogeneous clusters. This structured organization indicates that TrajVAE effectively learns a correlation between trajectories and latent vectors. Consequently, an optimization-based planning algorithm is likely to be effective.
@@ VI-C Brush Stroke Dynamics Modeling Experiments
The purpose of a stroke dynamics model is to differentiably render trajectories. We experiment with a variety of methods to do this and evaluate them both quantitatively and qualitatively in a controlled experiment. In total, we evaluated four methods:
@@ VI-C1 CNN
Our baseline, a convolutional neural network starting with a fully connected layer and followed by several transposed convolutions. This is analogous to FRIDA’s [10] renderer architecture, except it takes in full trajectory rather than the parameterization (length, bend, height) that FRIDA uses.